240 mL of 0.32 M aluminum chloride reacts with excess silver nitrate in a double displacement reaction. What mass of silver chloride precipitate is produced?


AlCl3 + 3 AgNO3 -> Al(NO3)3 + 3 AgCl

1
Expert's answer
2021-05-12T06:30:29-0400

The reaction is

"AlCl_3+3AgNO_3\Rightarrow Al{(NO_3)_2+3AgCl}"


The moles of Aluminium chloride =


"M=\dfrac{n_{AlCl_3}}{V(ml)}\times1000"


"0.32=\dfrac{n_{AlCl_3}}{240}\times1000"


"n_{AlCl_3}=0.077" moles


Now according to the symmetry of the equation:


1 mole of Aluminium chloride gave 3 moles of silver chloride

Hence 0.077 moles of Aluminium chloride gave "3\times0.077=0.231" moles of silver chloride.


Now mass of silver chloride =moles of silver chloride "\times" molar mass of silver chloride

="0.231\times143.32=33.10" gm


So 33.10 gm of silver chloride is precipitated in produced after the reaction.